An isosceles triangle has two sides of equal length. The length of the third side is 2 ft less than twice the length of one of the equal sides. Find the length of each side when the perimeter is 18 ft. (Enter your answers as a comma-separated list.)
Answer by rfer(16322)

x+x+2x-2=18
4x=18+2
x=20/4
x=5
2x-2=8
5, 5, 8